You are considering opening a copy service in the student union. You estimate your fixed cost at $ 15,000 and the variable cost of each copy sold at $ 0.01. You expect the selling price to average $ 0.06.

​a) Based on the given​ information, the​ break-even point in dollars​ = ​$ ? ​(round your response to the nearest whole​ number).

​b) Based on the given​ information, the​ break-even point in units​ = $ ? units ​(round your response to the nearest whole​ number).
Answer #1

Fixed cost(FC) = $15000

Variable cost(VC) = $0.01

Selling price(SP) = $0.06

a) Break even point in units = FC/(SP-VC) = 15000/(0.06-0.01) = 15000/0.05 = 300000 units

Break even point in dollars = Break even point in units x SP = 300000 x $0.06 = $18000

b) Break even point in units = FC/(SP-VC) = 15000/(0.06-0.01) = 15000/0.05 = 300000 Units